Suppose you invest R10 000 at a simple interest rate of 12% per year. What would be the value of the accumulated amount after 6 years


1
Expert's answer
2022-05-09T11:09:47-0400

"A=P(1+rt)"

"10000(1+0.12\cdot 6)=R\ 17200"